Projection Admin 3/19
To do this evening/ was completed: Weekly Update is available and staff screenings have been emailed out, scheduled on the LMS and blocked on the calendar.
Presentation Issues:
Preshow Clips/Playlist: Built the staff screenings playlists and scheduled them on the TMS.
The keys for Alto Knights open up tomorrow, so I'll get my eyes on it then.
QC'd Snow White, which sounds good at 5.0, perhaps we can raise it to 5.5 in the larger auditoriums, depending on how many patrons show, but for now I think it's perfectly well at 5.0.
Trailer Updates:
Special Events/ AV set up: Ran loops for Full Metal Jacket tonight. Should be a great show!
Trailer Updates:
Special Events/ AV set up: Ran loops for Full Metal Jacket tonight. Should be a great show!
REC finished transcoding in T1. Ingested this to the LMS and IMS so we can double check it tomorrow morning. Deleted the old version.
Ingested the CCAP/OCAP of Death of a Unicorn to the LMS.
The keys for They Cloned Tyrone open this evening, after 5:30pm. Please take a peek at it between the first and second round in T2 tonight, if you can. There will be 15minutes.
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): We received the CRU drive of Paper Moon this morning, which has been ingested to T1 and transferred to the LMS. I reached out to Deluxe for an eta on the key delivery and they're currently on hold. While I imagine they'll show up some time tomorrow or Friday, I've notified John just in case. Update: Received and ingested the keys.
Outgoing/Shipping (content): Uptown Girls will be on its way out today!
Print/video Inspection: Re-inspected, packaged up and printed out labels for Uptown Girls.
Asset tracking doc updated/checked?:
Maintenance/cleaning: Cleaned the projectors in T1.
Cue sheets updated/checked (every shift): Yes. Made a slight adjustment this morning to Paddington in T6 as it was an hour off.
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): We received the CRU drive of Paper Moon this morning, which has been ingested to T1 and transferred to the LMS. I reached out to Deluxe for an eta on the key delivery and they're currently on hold. While I imagine they'll show up some time tomorrow or Friday, I've notified John just in case. Update: Received and ingested the keys.
Received Zola via satellite. Ingested to the LMS. The keys open up Friday morning, so we can take a look then!
Outgoing/Shipping (content): Uptown Girls will be on its way out today!
Print/video Inspection: Re-inspected, packaged up and printed out labels for Uptown Girls.
Asset tracking doc updated/checked?:
Maintenance/cleaning: Cleaned the projectors in T1.
Cue sheets updated/checked (every shift): Yes. Made a slight adjustment this morning to Paddington in T6 as it was an hour off.
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Comments
Post a Comment